The average product of the 10 employees of ABC is 1000 bags. The firm is considering employing a Chinese worker whose marginal product is 900 bags. Each bag is sold at K 10.00

a) What is the average product after this worker is employed?

b) Should ABC employee the new Chinese worker if he demands a wage of K 8 000.00

c) What is the maximum wage ABC can pay the Chinese worker
Answer #1

A) Current AP = 1000

Current TP = AP.Q = 1000*10 = 10,000

MP = 900

NEW TP = 10,000+900 = 10,900

New AP = TP/Q = 10,900/11 = 990.91

B) Wage demanded = 8000

MRP of hiring additional worker = MP.P = 900*10 = 9000

Since MRP>W, it would be profitable to hire the worker.

Thus, the new worker should be hired

C) The maximum wage that can be offered would be equal to MRP = K 9000